OpenStudy (dinomancer):

If an object is dropped from a height of 200 feet, the function h(t)=-16t^2+200 gives the height of the object after t seconds. Approx. when will the object hit the ground? I would like an explanation on how to do this please.

OpenStudy (welshfella):

when it hits the ground h(t) equals 0. so you need to solve the equations -16t^2 + 200 -= 0 solve for t

OpenStudy (welshfella):

do you get that?

OpenStudy (dinomancer):

Did you mean to put the negative before the equal sign?

OpenStudy (anonymous):

kek

OpenStudy (dinomancer):

Still don't really understand this, if anyone could lend a hand that would be great.

OpenStudy (aveline):

Ok, here's the equation: h(t)=-16t^2+200 This means that after 0 seconds, the object is at 200 feet above the ground. The variable h(t) resembles the height above the ground. We want that to be 0. So, 0=-16t^2+200 First, subtract 200 from both sides -200=-16t^2 Next, divide by -16 12.5=t^2 Finally, square root √12.5=t The object will hit the ground after about 3.54 seconds.

OpenStudy (dinomancer):

Thanks, I totally understand this now :D

OpenStudy (welshfella):

No the negative was a typo